# LoongServe: Efficiently Serving Long-context Large Language Models with Elastic Sequence Parallelism

### v1

@@ 5.1. Dispatching
As for the performance gain, it is formulated as the impact on the input token latency of $R_{p,i}^{\prime}$:
@@ 5.5. Optimizations
Analytical Model based on SIB. The iteration time of the prefill phase under different scenarios guides the decision-making. Ideally, we can record them into SIB in advance and retrieve them at run time. However, there are innumerable combinations of requests with different input lengths executing at different DoPs. It is impossible to cover all the cases only by profiling. Therefore, we propose an analytical model to estimate them, which is formulated as follows:
@@
Other optimizations. LoongServe also employs other optimizations to improve the efficiency by filling small bubbles in the scheduling plan. Please refer to the Appendix A for more details.
@@
6. Implementation
LoongServe is implemented in approximately 15K lines of code based on C++, CUDA, Python, and OpenAI Triton (Tillet et al., 2019), and reuses some components from LightLLM (Team, 2023) and vLLM (Kwon et al., 2023). The front end of LoongServe is the same as LightLLM and is similar to OpenAI API (OpenAI, 2023). Users can send requests to LoongServe without modifications if they have used LightLLM previously.
For each elastic instance, it manages its corresponding key-value cache pool by using PagedAttention (Kwon et al., 2023) at the granularity of a single token. Although StripedAttention performs pretty well in long sequences, it causes redundant computation in short sequences due to its special causal attention mask. We tune the tile size to skip most redundant computations in short sequences. We also reduce the lifecycle of extra parameters introduced by StripedAttention in shared memory to improve the occupancy of streaming multiprocessors (SMs). For the decoding phase, we implemented a customized version of Flash-Decoding (Dao, 2023) with extra parameters to support ESP. All the above optimizations are compatible with MHA, MQA, and GQA, and have the same accuracy as the original implementations.
@@ 7. Evaluation
7.1. Experimental Setup
Model. We use the LWM-1M-Text model (Liu et al., 2024) as the long-context LLM model in our evaluation. It is the open-sourced pre-trained LLM with the largest context window size (1M tokens) when we started experiments. Besides, it uses the same model architecture as Llama-2-7B (Touvron et al., 2023b), which is widely used in practice.
@@ 7.1. Experimental Setup
Workloads. Similar to prior work (Li et al., 2023d; Kwon et al., 2023; Zhong et al., 2024), the arrival pattern of requests is generated by a Poisson process. The input lengths and output lengths of requests are sampled from the following real-world datasets:
@@
  • LV-Eval (Yuan et al., 2024): It is the dataset containing the longest requests when we started experiments. It mainly contains many long-context question-answer tasks. When converting words into tokens, the range of sequence length in this dataset is 15.1K - 497.3K tokens.
@@
  • DeepSpeed-MII (dee, 2024)222commit hash: 773b735d6294a98dd842d82ef024d0d9b050f66aa : It proposes Dynamic SplitFuse (Holmes et al., 2024) to decompose long input sequences into small chunks to prevent performance degradation of the decoding phase when serving both phases. However, when serving requests with long context larger than 32K tokens, we encounter the ”illegal memory access” error, so we only evaluate it on ShareGPT. The tensor parallelism is set to 8.
  • LightLLM w/ SplitFuse (Team, 2023)333commit hash: e2b5168a50f24d960ead314b0649428e35381f80 : To evaluate the performance of SplitFuse for long sequences, we also evaluate SplitFuse provided by LightLLM. Similar to SARATHI (Agrawal et al., 2023), it needs to set the chunk size. We set the chunk size as the ideal ”P:D ratio” described in SARATHI (Agrawal et al., 2023) by calculating it for each dataset before experiments, although it is unknown in practice. The tensor parallelism is set to 8.
  • DistServe (Zhong et al., 2024): It proposes prefill-decode disaggregation to reduce the impact of the prefill phase on the decoding phase. We contact the authors to get the source code and set parallelism strategies for it. Because its parallelism strategies are restricted by the head of the model and requests with long context need large GPU memory, we use four GPUs for the prefill phase and four GPUs for the decoding phase. The DoP for each phase is set to 4.
@@
Per-token latency.    Input latency.    Output latency.
Figure 11. Multi-node Performance.
Metrics. We focus on the serving throughput. For each request rate, we measure the normalized per-token latency, i.e., the mean of requests’ end-to-end latency divided by their sequence lengths, normalized input latency, i.e., the mean of prefill phase time divided by the input lengths, and the normalized output latency, i.e., the mean of decoding phase time divided by the output lengths, for each system. To compare different systems, we set a latency service level objective (SLO) and compare the maximum throughput under this SLO. Similar to prior work (Wu et al., 2023), we set SLO to 25$\times$ of the latency under the light load.
7.2. End-to-End Performance
We compare LoongServe with four baselines on four real-world workloads in three key metrics. Figure 10 shows the results. Because LoongServe often uses different elastic instances to execute different phases, the decoding phase is well protected from the impact of the prefill phase. As a result, the output latency of LoongServe achieves low latency and is significantly better than other baselines (nearly infinite output throughput improvement). For the prefill phase, LoongServe accelerates them by setting the appropriate DoP to avoid blocking pending prefill phases. In contrast, vLLM consistently treats different requests in the same way. On the ShareGPT, vLLM wastes resources on short requests with poor scalability. On other datasets, it causes severe interference between the decoding phase and the prefill phase. LoongServe improves the total throughput and input throughput by up to 4.64$\times$ and 4.00$\times$. DeepSpeed MII and LightLLM w/ SplitFuse try to split the input sequence into chunks to protect the decoding phase. However, decomposing the input sequence makes the prefill phase inefficient. Furthermore, for long sequences in L-Eval and LV-Eval, the prefill phase still causes severe interference with the decoding phase, because the ”P:D” ratio is high. Compared to them, LoongServe improves the total throughput and input throughput by up to 3.85$\times$ and 3.37$\times$. DistServe disaggregates the prefill phase and the decoding phase to reduce the interference between them. However, it means that each phase can only use half of the GPUs (four GPUs). On the ShareGPT, four GPUs are not enough to serve lots of requests in the decoding phase. On the L-Eval, four GPUs are not enough to serve lots of requests in the prefill phase. On the LV-Eval and Mixed, four GPUs for each phase even do not have enough memory to serve requests with long context and thus trigger the Out-of-Memory (OOM) error, so it is not shown in these two rows of figures. Compared to DistServe, LoongServe improves the total throughput and input throughput by up to 5.81$\times$ and 3.58$\times$.
@@ 7.3. Multi-Node Performance
We also evaluate the multi-node performance of LoongServe on a 16 GPU cluster. In this experiment, we deploy baselines in each server and use the same parallelism strategies as the single-node evaluation. LoongServe also uses the same model parallelism strategy and extends the ESP to 8. We also use the same Mixed dataset. Figure 11 shows the results. In the multi-node setting, LoongServe scales well and achieves the best performance in all metrics by setting the appropriate DoP for each request. Therefore, LoongServe avoids unnecessary communication overhead for short requests and enlarges the parallelism for long requests. As a result, LoongServe improves the total throughput and input throughput by up to 1.86$\times$ and 1.72$\times$ compared to vLLM, 3.37$\times$ and 3.11$\times$ compared to LightLLM w/ SplitFuse, while significantly reducing the output latency under all request rates.
@@ 7.4. Ablation Study
7.4.1. Effectiveness of Elastic Sequence Parallelism
To show the effectiveness of elastic sequence parallelism, we conduct an ablation study on the P90 goodput (throughput of requests under the SLO) (Zhong et al., 2024) of different parallelism strategies under different sequence length distributions. Baselines include traditional tensor parallelism, i.e., LoongServe w/o ESP (TP=8), static hybrid parallelism, i.e., LoongServe w/o ESP (TP=2, SP=4), and parallelism with replication, i.e., LoongServe w/o ESP (TP=2) $\times$ 4. To simulate different scenarios, we sample the sequence lengths from the Mixed dataset with different Zipf parameters. Because the baseline with replication is not able to serve requests with long context, we limit the maximum length of requests to 200K tokens.
Figure 12 shows the results. As shown in the figure, only introducing sequence parallelism is not enough to achieve good performance. Neither static hybrid parallelism nor parallelism with replication handles the dynamic inference workload. In contrast, LoongServe dynamically adjusts strategies for the dynamic workload and consistently achieves P90 goodput improvement by 2.33$\times$, 1.98$\times$, and 1.53$\times$ under different sequence distributions. P90 goodput also indicates that ESP is beneficial for most requests.
@@ 7.4.1. Effectiveness of Elastic Sequence Parallelism
7.4.2. Scaling Overhead
We evaluate the overhead of scaling down and scaling up under different batch sizes and input lengths by timing the time to forward a batch with and without scaling, respectively. Figure 13 shows the results. For scaling down, we only introduce a negligible overhead (less than 2%) for all batch sizes and prompt lengths. For scaling up, on large batch sizes, since the compute-memory ratio is high for matrix multiplication operations (including projection for query, key, and value, as well as matrix multiplication for the FFN), distributing the computation to more instances can significantly reduce the computation time, leading to a 2$\times$ improvement in per-iteration latency. On small batch sizes, the overhead of scaling up is higher because of the overhead introduced by communication and synchronization. However, the overhead is still acceptable (less than 10%) and LoongServe dynamically uses the best one.
7.4.3. Accuracy of Analytical Model
At last, we evaluate the accuracy of analytical models in different parallelism strategies. As shown in Figure 14, the analytical model of LoongServe achieves high accuracy (less than 10% deviation) for different batches of requests with different sequence lengths in different parallelism strategies. It is reliable to guide the LoongServe global manager.
@@ 8. Related Work
Existing LLM Serving Systems. Existing LLM serving systems adopt efficient GPU operators for the long sequence requests, such as Flash Attention (Dao et al., 2022) and Flash-Decoding (Dao, 2023). They are orthogonal to our work and are integrated into LoongServe. To mitigate the impact of the long context, SARATHI (Agrawal et al., 2023) and DeepSpeed-FastGen (Holmes et al., 2024) split the long context into chunks and process them chunks by chunks, but they still incur interference between two phases (Zhong et al., 2024). SplitWise (Patel et al., 2023), DistServe (Zhong et al., 2024), and TetriInfer (Hu et al., 2024) disaggregate two phases into different groups of GPUs to avoid interference, but their static parallelism and partition strategies are not flexible to handle the dynamic workload. Infinite-LLM (Lin et al., 2024) tries to alleviate the GPU fragmentation across instances. However, it still needs periodic key-value tensor migration to maintain the locality and does not consider elastic resource demand between different requests or different phases. LoongServe proposes a novel elastic scaling mechanism without additional overhead and ESP to handle dynamic workload based on requests’ resource demand without locality constraints.
@@
Sequence Parallelism. Many works (Liu et al., 2023; Brandon et al., 2023; Jacobs et al., 2023; Li et al., 2023a; Korthikanti et al., 2023; Li et al., 2023c) are proposed to accelerate long-context LLM training. One class of work still uses TP for the attention layers and suffers from the same problems as TP. Another class of work like Striped Attention (Brandon et al., 2023) parallelizes the attention mechanism in the sequence dimension. All these works focus on LLM training and their DoP are fixed. Our work targets the serving scenario, supporting the decoding phase, dynamic DoP, and efficient key-value cache management.
Long-context LLM with Accuracy Loss. Another type of Long-context LLM trades off accuracy for efficiency. Some works (Child et al., 2019; Pagliardini et al., 2023; Sun et al., 2024) change the attention mechanism to reduce computation. Recent works (Zhang et al., 2024; Xiao et al., 2024; Adnan et al., 2024) also try to prune the key-value cache to reduce the memory footprint. All these works incur accuracy loss and are not widely used. LoongServe does not affect the accuracy of the original LLM and provides a more efficient serving mechanism. Besides, LoongServe is compatible with MQA (Shazeer, 2019), GQA (Ainslie et al., 2023), and MoE (Jiang et al., 2024) to reduce the memory footprint and computational complexity.
Elastic Training. Many works focus on elastic training for deep neural networks (DNNs) (Li et al., 2023b; Gu et al., 2023; Ma et al., 2021; Gu et al., 2019; Qiao et al., 2021; Hwang et al., 2021; Peng et al., 2018; Jayaram Subramanya et al., 2023). Compared to them, LoongServe proposes a new parallelism strategy, ESP, rather than using data parallelism as in the training phase. LoongServe also considers more unique factors in the LLM serving, such as key-value tensor management and request batching, under more strict latency constraints.

### v2

@@ 5.1. Dispatching
In this equation, the output length of requests is the number of existing output tokens. This estimation is similar to the output length-based waiting time estimation in previous LLM serving works (Wu et al., 2024; Liu et al., 2024a; Patke et al., 2024). As for the performance gain, it is formulated as the impact on the input token latency of $R_{p,i}^{\prime}$:
@@ 5.5. Optimizations
Analytical model based on SIB. The iteration time of the prefill phase under different scenarios guides the decision-making. Ideally, we can record them into SIB in advance and retrieve them at run time. However, there are massive combinations of requests with different input lengths executing at different DoPs. It is impossible to cover all the cases only by profiling. Therefore, we propose an analytical model to estimate them, which is formulated as follows:
@@
6. Implementation
LoongServe is implemented in approximately 15K lines of code based on C++, CUDA, Python, and Triton (Tillet et al., 2019), and reuses some components from vLLM (Kwon et al., 2023) and LightLLM (Team, 2023). The front end of LoongServe is similar to OpenAI API (Achiam et al., 2023). Users send requests to LoongServe based on the front-end API. The code of LoongServe is open-source and is publicly available at https://github.com/LoongServe/LoongServe.
@@ 6. Implementation
For each elastic instance, it manages its corresponding key-value cache pool by using PagedAttention (Kwon et al., 2023) at the granularity of a single token. Although StripedAttention performs pretty well on long sequences, it causes redundant computation in short sequences due to its special causal attention mask. We tune the tile size to skip most redundant computations in short sequences. We also reduce the lifecycle of extra parameters introduced by StripedAttention in shared memory to improve the occupancy of streaming multiprocessors (SMs). For the decoding phase, we implement a custom version of Flash-Decoding (Dao, 2023) with extra parameters to support ESP. All the above optimizations are compatible with MHA, MQA, and GQA, and have the same accuracy as the original implementations.
@@ 7. Evaluation
7.1. Methodology
Model. We use the LWM-1M-Text model (Liu et al., 2024b) as the long-context LLM model in our evaluation. It is an open-source pre-trained LLM with the largest context window size (1 million tokens) when we started our evaluation. Besides, it uses the same model architecture as Llama-2-7B (Touvron et al., 2023b), which is widely used in practice.
@@ 7.1. Methodology
Workloads. Similar to prior work (Li et al., 2023d; Kwon et al., 2023; Zhong et al., 2024), the arrival pattern of requests is generated by a Poisson process. The input lengths and output lengths of requests are sampled from the following real-world datasets.
@@
  • LV-Eval (Yuan et al., 2024): It is the dataset containing the longest requests when we started our evaluation. It contains long-context question-answer tasks. When converting words into tokens, the range of sequence length in this dataset is 15.1K - 497.3K tokens.
@@
Per-token latency.         Input latency.         Output latency.
Figure 11. Multi-node performance.
@@
  • DeepSpeed-MII (dee, 2024) [^2] DeepSpeed MII, https://github.com/microsoft/DeepSpeed-MII, commit hash: 773b735d6294a98dd842d82ef024d0d9b050f66aa : It proposes Dynamic SplitFuse (Holmes et al., 2024) to decompose long input sequences into small chunks to prevent performance degradation of the decoding phase when serving both phases. However, when serving requests with long context larger than 32K tokens, we encounter the ”illegal memory access” error, so we only evaluate it on ShareGPT. The tensor parallelism is set to 8.
  • LightLLM w/ SplitFuse (Team, 2023) [^3] DistServe, https://github.com/LLMServe/DistServe, commit hash: e2b5168a50f24d960ead314b0649428e35381f80 : To evaluate the performance of SplitFuse for long sequences, we also evaluate SplitFuse provided by LightLLM. Similar to SARATHI (Agrawal et al., 2023), it needs to set the chunk size. We set the chunk size as the ideal ”P:D ratio” described in SARATHI (Agrawal et al., 2023) by calculating it for each dataset before experiments, although it is unknown in practice. The tensor parallelism is set to 8.
  • DistServe (Zhong et al., 2024): It proposes prefill-decode disaggregation to reduce the impact of the prefill phase on the decoding phase. We contact the authors to get the source code and set parallelism strategies for it. Because its parallelism strategies are restricted by the head of the model and requests with long context need large GPU memory, we use four GPUs for the prefill phase and four GPUs for the decoding phase. The DoP for each phase is set to 4, which is the best parallelism strategy searched by DistServe’s simulator and is validated on the testbed.
@@
Metrics. We focus on the serving throughput. For each request rate, we measure the normalized per-token latency, i.e., the mean of requests’ end-to-end latency divided by their sequence lengths, normalized input latency, i.e., the mean of prefill phase time divided by the input lengths, and the normalized output latency, i.e., the mean of decoding phase time divided by the output lengths, for each system. To compare different systems, we set a latency service level objective (SLO) and compare the maximum throughput under this SLO. Similar to prior work (Wu et al., 2023; Li et al., 2023d; Kwon et al., 2023), we set SLO to 25$\times$ of the inference latency.
@@
7.2. End-to-End Performance
We compare LoongServe with four baselines on four real-world workloads in three key metrics. Figure 10 shows the results. Because LoongServe uses different elastic instances to execute different phases, the decoding phase is well protected from the impact of the prefill phase. As a result, the output latency of LoongServe achieves low latency and is significantly better than other baselines. For the prefill phase, LoongServe accelerates them by setting an appropriate DoP to avoid blocking pending prefill phases. In contrast, vLLM consistently treats different requests in the same way. On the ShareGPT, vLLM wastes resources on short requests with poor scalability. On other datasets, it causes severe interference between the decoding phase and the prefill phase. LoongServe improves the total throughput and input throughput by up to 4.64$\times$ and 4.00$\times$. DeepSpeed MII and LightLLM w/ SplitFuse try to split the input sequence into chunks to protect the decoding phase. However, decomposing the input sequence makes the prefill phase inefficient. Furthermore, for long sequences in L-Eval and LV-Eval, the prefill phase still causes severe interference with the decoding phase, because the ”P:D ratio” (prefill/decoding ratio) (Agrawal et al., 2023) is high. Compared to them, LoongServe improves the total throughput and input throughput by up to 3.85$\times$ and 3.37$\times$. DistServe disaggregates the prefill phase and the decoding phase to reduce the interference between them. However, due to the isolation between the two phases, it means that each phase can only use half of the GPUs (four GPUs) and the longest length of a request is bounded by the minimum GPU capacity of two phases, i.e., GPU memory capacity of four GPUs. On the ShareGPT, four GPUs are not enough to serve lots of requests in the decoding phase. On the L-Eval, four GPUs are not enough to serve lots of requests in the prefill phase. On the LV-Eval and Mixed, four GPUs for each phase even do not have enough memory to serve requests with long context and thus trigger the Out-of-Memory (OOM) error, so it is not shown in these two rows of figures. In contrast, with an unified distributed KV cache pool, LoongServe can serve them well and achieve the best performance. Compared to DistServe, LoongServe improves the total throughput and input throughput by up to 5.81$\times$ and 3.58$\times$.
(a) Effectiveness of elastic scale-up.
(b) Frequency of elastic scale-up.
Figure 13. Ablation study of elastic scale-up.
@@ 7.3. Multi-Node Performance
We also evaluate the multi-node performance of LoongServe on a 16-GPU cluster. In this experiment, we deploy baselines on each server and use the same parallelism strategies as the single-node evaluation. LoongServe also uses the same model parallelism strategy and extends the ESP to 8. We also use the same Mixed dataset. Figure 11 shows the results. In the multi-node setting, LoongServe scales well and achieves the best performance in all metrics by setting an appropriate DoP for each request. Therefore, LoongServe avoids unnecessary communication overhead for short requests and enlarges the parallelism for long requests. As a result, LoongServe improves the total throughput and input throughput by up to 1.86$\times$ and 1.72$\times$ compared to vLLM, 3.37$\times$ and 3.11$\times$ compared to LightLLM w/ SplitFuse, while significantly reducing the output latency under all request rates. The results show that LoongServe can effectively handle the dynamic workload and achieve high performance in the multi-node setting because it can choose appropriate DoPs for diverse requests at different phases.
@@ 7.4. Ablation Study
Effectiveness of elastic sequence parallelism. To show the effectiveness of elastic sequence parallelism, we conduct an ablation study on the P90 goodput (throughput of requests under the SLO) (Zhong et al., 2024) of different parallelism strategies under different sequence length distributions. Baselines include traditional tensor parallelism, i.e., LoongServe w/o ESP (TP=8), static hybrid parallelism, i.e., LoongServe w/o ESP (TP=2, SP=4), and parallelism with replication, i.e., LoongServe w/o ESP (TP=2) $\times$ 4. To simulate different scenarios, we sample the sequence lengths from the Mixed dataset with different Zipf parameters. Because the baseline with replication is not able to serve requests with long context, we limit the maximum length of requests to 200K tokens.
Figure 12 shows the results. As shown in the figure, only introducing sequence parallelism is not enough to achieve good performance. Neither static hybrid parallelism nor parallelism with replication handles the dynamic inference workload. In contrast, the scalable four-step scheduling algorithm of LoongServe dynamically adjusts strategies for the dynamic workload and achieves P90 goodput improvement by 2.33$\times$, 1.98$\times$, and 1.53$\times$ under different sequence distributions. P90 goodput also indicates that ESP is beneficial for most requests.
Elastic scale-up. We further evaluate the effectiveness of elastic scale-up. Similar to Figure 12, we first conduct an ablation study on the P90 goodput between LoongServe with and without elastic scale-up. As shown in Figure 13a, when serving the ShareGPT dataset, the P90 goodput of LoongServe with elastic scale-up is 2.87$\times$ higher than LoongServe without elastic scale-up. The reason is that requests from the ShareGPT dataset have a relatively short input length and long output length, which requires frequent scaling up as the output length continuously increases. As shown in Figure 13, we measure the frequency of triggering elastic scale-up when the request arrival rate is 25 requests per second on the ShareGPT dataset. Each bin in the figure represents the number of elastic scale-up operations triggered by the global manager in a 10-second interval. On average, the global manager triggers 7.12 elastic scale-up operations per 10 seconds, which indicates the necessity of elastic scale-up to handle the dynamic workload.
@@
Scaling overhead. We evaluate the overhead of scaling down and scaling up under different batch sizes and input lengths by timing the time to forward a batch with and without scaling, respectively. Figure 14 shows the results. For scaling down, we only introduce a negligible overhead (less than 2%) for all batch sizes and prompt lengths. For scaling up, on large batch sizes, since the compute-memory ratio is high for matrix multiplication operations (including projection for query, key, and value, as well as matrix multiplication for the FFN), distributing the computation to more instances can significantly reduce the computation time, leading to a 2$\times$ improvement in per-iteration latency. On small batch sizes, the overhead of scaling up is higher because of the overhead introduced by communication and synchronization. However, the overhead is still acceptable (less than 10%) and LoongServe dynamically uses the best one.
Accuracy of analytical model. At last, we evaluate the accuracy of analytical models in different parallelism strategies. As shown in Figure 15, the analytical model of LoongServe achieves high accuracy (less than 10% deviation) for different batches of requests with different sequence lengths in different parallelism strategies. It is reliable to guide the LoongServe global manager.
@@ 8. Related Work
LLM serving systems. Existing LLM serving systems adopt efficient GPU kernels for long sequence requests, such as Flash Attention (Dao et al., 2022) and Flash-Decoding (Dao, 2023). They are orthogonal to our work and are integrated into LoongServe. To mitigate the impact of the long context, SARATHI (Agrawal et al., 2023) and DeepSpeed-FastGen (Holmes et al., 2024) split the long context into chunks and process them chunks by chunks, but they still incur interference between two phases (Zhong et al., 2024). SplitWise (Patel et al., 2023), DistServe (Zhong et al., 2024), and TetriInfer (Hu et al., 2024) disaggregate two phases into different groups of GPUs to avoid interference, but their static parallelism and partition strategies are not flexible to handle the dynamic workload. Infinite-LLM (Lin et al., 2024) tries to alleviate the GPU fragmentation across instances. However, it still needs periodic key-value tensor migration to maintain the locality and does not consider elastic resource demand between different requests or different phases. LoongServe proposes a novel elastic scaling mechanism without additional overhead and ESP to handle dynamic workload based on requests’ resource demand without locality constraints.
Sequence parallelism. Many works (Liu et al., 2023; Brandon et al., 2023; Jacobs et al., 2023; Li et al., 2023a; Korthikanti et al., 2023; Li et al., 2023c) are proposed to accelerate long-context LLM training and have demonstrated the benefit of sequence parallelism under both single-host and multi-host settings. One class of work still uses TP for the attention layers and suffers from the same problems as TP. Another class of work like Striped Attention (Brandon et al., 2023) parallelizes the attention mechanism in the sequence dimension. All these works focus on LLM training and their DoP are fixed. Our work targets the serving scenario, supporting the decoding phase, dynamic DoP, and efficient key-value cache management.
Long-context LLM with accuracy loss. Another type of Long-context LLM trades off accuracy for efficiency. Some works (Child et al., 2019; Pagliardini et al., 2023; Sun et al., 2024) change the attention mechanism to reduce computation. Recent works (Zhang et al., 2024; Xiao et al., 2024; Adnan et al., 2024) also try to prune the key-value cache to reduce the memory footprint. All these works incur accuracy loss. LoongServe does not affect the accuracy of the original LLM and provides a more efficient serving mechanism. Besides, LoongServe is compatible with MQA (Shazeer, 2019), GQA (Ainslie et al., 2023), and MoE (Jiang et al., 2024) to reduce the memory footprint and computational complexity.
Elastic training. Many works focus on elastic training for deep neural networks (DNNs) (Li et al., 2023b; Gu et al., 2023; Ma et al., 2021; Gu et al., 2019; Qiao et al., 2021; Hwang et al., 2021; Peng et al., 2018; Jayaram Subramanya et al., 2023). Compared to them, LoongServe proposes a new parallelism strategy, elastic sequence parallelism, rather than using data parallelism as in the training phase. LoongServe also considers more unique factors in the LLM serving, such as decoding phase, key-value tensor management and request batching, under more strict latency constraints.